An agent sells life insurance policies to five equally aged, healthy people. According to recent data, the probability of a person living in these conditions for 30 years or more is 2/3. Calculate the probability that after 30 years:

a. All five people are still living.

b. At least three people are still living.

c. Exactly two people are still living.


1
Expert's answer
2021-05-10T18:28:56-0400

Let us consider person living in these conditions for "30" years or more is a success.

Then probability of success is "\frac{2}{3}" and probability of failure is "\frac{1}{3}".

"\therefore p=\frac{2}{3}" and "q=\frac{1}{3}".

Now we know that for binomial distribution probability of getting "r" success from "n" trial is "={^n}C_r.p^r.q^{(n-r)}"

(a) Here "n=5,r=5"

"\therefore" Probability of all five people are still living is "={^5}C_5.(\frac{2}{3})^5.(\frac{1}{3})^{(5-5)}=(\frac{2}{3})^5"

(b) Here "n=5" .We have to calculate the probability for "r=3,4,5" .

"\therefore" Probability of at least three people are still living is "={^5}C_3.(\frac{2}{3})^3.(\frac{1}{3})^{(5-3)}+{^5}C_4.(\frac{2}{3})^4.(\frac{1}{3})^{(5-4)}+{^5}C_5.(\frac{2}{3})^5.(\frac{1}{3})^{(5-5)}"

"=24.(\frac{2^3}{3^5})"

(c) Here "n=5,r=2"

Probability of exactly two people are still living "={^5}C_2.(\frac{2}{3})^2.(\frac{1}{3})^{(5-2)}" "={^5}C_5.(\frac{2}{3})^2.(\frac{1}{3})^{3}" "=\frac{4}{3^5}"